1 16.2 P011 My Ask Your An electron is at the origin. (a) Calculate the electric potential VA at point A,x- 0.310 om. (b) Calculate the electric potential va at point B, x . o.715 cm. (c) Would a negatively charged particle placed at point A necessarily go through this same potential dilference upon reaching point B? Explain This answer has not been graded yet Need Help? RedN
Answer #1

a)

Electric potential at point a can be written as

Va= q / 4π epsilonr

Where

epsilon= 8.85 10 12 = permittivity of the free space

r= distance from the Origin to the x

r=x= 0.31cm=0.31 10 2m

Therefore

Va 1.6 10 19/4 3.14 8.85 10 120.31 102

Va= 0.464*10 6 V

Where charge of electron q= 1.6*10 19 C

b)

Now electric potential Vb at point x= 0.715cm

Vb = q/4πepsilonr

Where r=x= 0.715*10 2m

Vb- 1.610 19/43.14 * 8.85 10 12 0.71510 2

Vb= 0.210 * 10 6 V

c)

The potential difference Vb - Va is

Vb - Va= (0.210 - 0.464)* 10 6

Vb - Va= -0.254 V